I dont know how to write this but I will try, I am struggling to get pass my algebra class, The question I have is
x+1=9x^3+9x^2
also I have a problem I dont know how to type it but I will try
it has the radical equation sign I think its called it look like the old division like a check mark so it is that under it is x+10 than =x-2. Could you please help me
Thank You
Debbie

Answer by stanbon(75887)

x+1=9x^3+9x^2
9x^3+9x^2-x-1 = 0
Factor:
9x^2(x+1)-(x+1) = 0
(x+1)(9x^2-1) = 0
(x+1)(3x-1)(3x+1) = 0
x = -1 or x = 1/3 or x = -1/3
===========================================
also I have a problem I dont know how to type it but I will try
it has the radical equation sign I think its called it look like the old division like a check mark so it is that under it is x+10 than =x-2.
-----
sqrt(x+10) = x-2
----
Square both sides to get:
x+10 = x^2 -4x + 4
Simplify:
x^2 - 5x -6 = 0
Factor:
(x-6)(x+1) = 0
x = 6 or x = -1
=======================
Cheers,
Stan H.